OpenStudy (anonymous):

NEED HELP QUICKLY PLEASE, WILL GIVE MEDAL. The table below shows the squares of different numbers: Number (x) 2 -2 3 -3 Square of the number (y) 4 4 9 9 Part A: Does the table represent y as a function of x? Justify your answer. Part B: The total cost f(x), in dollars, for renting a rowboat for x hours is shown below: f(x) = 20 + 3x What is the value of f(150), and what does f(150) represent?

OpenStudy (anonymous):

@dtan5457

OpenStudy (triciaal):

part a yes y = f(x) x independent for any value of x only one value for y y = x^2

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

do you know part B and C?

OpenStudy (triciaal):

part B is not clear and I don't see a part C

OpenStudy (triciaal):

first says f(x) the total cost then its asking the value of 150

OpenStudy (anonymous):

okay. Part B we need to solve the Equation F(x) = 20 + 3x

OpenStudy (triciaal):

so 150 = 20 + 3x to find x -20 from each side then divide by 3 this gives the number of hours

OpenStudy (anonymous):

21?

OpenStudy (triciaal):

if the total paid was 150 but if 150 is the number of hours then x = 150 and the amount paid will be 20 + 3*150

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

470

OpenStudy (triciaal):

what does Part C ask?

OpenStudy (anonymous):

I just realized there is not part C I thought there was but it was just a sub section of part B, thank you for your help.

OpenStudy (anonymous):

do you have time for another question?

OpenStudy (triciaal):

do you understand the difference between f(x) = 150 and f(150)?

OpenStudy (triciaal):

open a new post with the other ?

OpenStudy (anonymous):

okay

OpenStudy (triciaal):

f(x) = 150 vs f(150) total cost paid vs 150 hours